OpenStudy (anonymous):

How many carbon atoms are there in 52.06 g of carbon dioxide? show math please!

OpenStudy (anonymous):

Molar mass CO2 = 44.0096g/mol The 52.06g = 52.06/44.0096 = 1.1829 mol CO2 1mol CO2 = 6.022*10^23 molecules 1.1829mol = 1.1829*6.022*10^23 = 7.124*10^23 CO2 molecules 1molecule CO2 contains 1 carbon atom Therefore there are 7.124*10^23 carbon atoms.